A store sells walnuts for $ 3.00 per pound and pecans for $ 5.00 per pound. How many pounds of each should be used to make a mixture that weighs 10 pounds and costs $ 3.70 per pound?
Answer by lwsshak3(11628)

A store sells walnuts for $ 3.00 per pound and pecans for $ 5.00 per pound. How many pounds of each should be used to make a mixture that weighs 10 pounds and costs $ 3.70 per pound?
**
let x= lbs of walnut in mixture
10-x=lbs of pecans in mixture
..
3x+5(10-x)=10*3.70
3x+50-5x=37
2x=13
x=13/2=6.5 lbs
10-x=3.5 lbs
ans:
lbs of walnut in mixture=6.5
lbs of pecans in mixture=3.5